OpenStudy (anonymous):

2x=y+11 3y+7=-(1/2)x

OpenStudy (anonymous):

y -2x = -11 3y+1/2 (x) = -7 x3 equation no.1 : 3y - 6x = -33 solve : -6x - 1/2 x = -33 -6.5x = - 33 x = 33 / 6.5 = 5,08 y = -11 - 2 ( 5.08 ) = -21.16 I guess that its -1/2(x) for your 2nd equation ?